OpenStudy (anonymous):

2 cars one going 10 miles/hr faster than the other equal distance apart 168 miles and meet in 1.6 hrs what is the rate the cars are driving?

OpenStudy (anonymous):

first car speed = x secon car speed = x+10 s=d/t t=d/s t by first car =168/x second car= 168/(x+10) t is 1.6 hours=>1.6 hr 1.6=168/x x=168/1.6=105 miles /hr => speed fpr the first car 115 speed of the second car

OpenStudy (anonymous):

agree

OpenStudy (anonymous):

medal de

OpenStudy (anonymous):

medal de

OpenStudy (saifoo.khan):

narayain's right^

OpenStudy (anonymous):

WRONG - 58 mph and 48 mph

OpenStudy (anonymous):

Thx anyway